QUESTION: SOLVE EACH OF THE FOLLOWING SYSTEMS BY SUBSTITUTION
5X-2Y=-5
Y-5X=3 I believe your substituting for the y

Answer by checkley71(8403)

Y-5X=3
Y=5X+3
THUS
5X-2(5X+3)=-5
5X-10X-6=-5
-5X=-5+6
-5X=1
X=1/-5
X=-1/5
Y-5*-1/5=3
Y+1=3
Y=3-1
Y=2